# Grasp2Grasp: Vision-Based Dexterous Grasp Translation via Schrödinger Bridges

### v1

@@ 5.1 Experimental Setup
Method Success Rate (%$\uparrow$ Diversity (rad)$\uparrow$ 6D GWH IoU (%$\uparrow$
@@
Metrics. We assess performance using several quantitative metrics. First, we report the grasp success rate using IsaacGym [53], following the GenDexGrasp [42] evaluation protocol. A grasp is considered successful if it maintains a stable hold on the object under six perturbation trials, where external forces are applied along ±$\pm x$, ±$\pm y$, and ±$\pm z$ axes. We apply each force as a uniform acceleration of $0.5m/s^{2}$ for 60 simulation steps and measure whether the object translates more than 2 cm. A grasp passes if it withstands all six trials. Simulation parameters include a friction coefficient of 10 and an object density of 10,000. We use IsaacGym’s built-in positional controller to achieve the desired joint configurations. In addition to the success rate, we report the diversity of the generated grasps, computed as the average standard deviation across all revolute joints, as well as the functional similarity between source and translated grasps. The latter is measured using the full 6D IoU of GWH as defined in Eq. 11.
Baselines. We evaluate all four variants of our model, each trained using a distinct physics-based OT cost introduced in Sec 4.2. As baselines, we compare against RobotFingerPrint [37], which is a recent grasp transfer method using a unified gripper coordinate space, and a diffusion-based model trained with Contact Map Similarity OT, conditioned on the source contact map.
@@ 5.2 Quality of Translated Grasps
Among all methods, our approach using the Jacobian-based OT cost achieves the highest mean success rate and outperforms other baselines by a substantial margin. Notably, the Jacobian variant also shows consistent improvements across all tasks, demonstrating its ability to preserve functionally effective contact configurations despite differences in morphology. The GWH-based model also performs competitively, which suggests that both wrench- and control-aligned transport costs are particularly effective for grasp transfer. Qualitative results in Fig. 3 (left) also show that our method can generate stable grasps across variants, even when the source grasp is sub-optimal. In contrast, object-agnostic methods like RFP may fail to form valid grasps due to a lack of fine-grained contact reasoning.
Figure 3: (left) Qualitative results. Our method generates stable and consistent grasps across different OT variants, even when the source grasp is sub-optimal where RFP [37] fails. (right) Failure Modes where our method struggles with thin-shell objects due to challenging geometry.
On the diversity metric, the Contact OT model achieves the highest average joint-space variation. This highlights that our methods can generate diverse grasp configurations while maintaining physical feasibility. Interestingly, other OT costs show lower diversity despite achieving high success rates, indicating that geometric alignment alone may lead to more conservative grasps.
@@ 6 Discussion
Limitations. Despite its strengths, our approach has some limitations. Most notably, as shown in Fig. 3 (right), the performance drops in scenarios involving thin-shell objects with ambiguous or minimal contact regions, where the valid grasp configuration space is largely confined. Additionally, we observe that the results for the Shadow hand are consistently lower than for the Allegro hand across multiple metrics, which aligns with prior findings in the MultiGripperGrasp dataset [11] due to its more complex kinematics and larger workspace. These discrepancies suggest that improving dataset quality and coverage for specific hands may further enhance generalization and translation fidelity.
@@
Appendix A Implementation Details of the Overall Objectives
In our method, we adopt the [SF]2[78] training framework to jointly learn the score function $s_{\theta}(t,x)$ and the drift field $v_{\theta}(t,x)$ of the corresponding SDE and its probability flow ODE. We follow the standard formulation of the [SF]2M loss:
$\mathcal{L}_{\mathrm{[SF]}^{2}\mathrm{M}}(\theta)=\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}t\sim\mathcal{U}(0,1)\\ x\sim p_{t}(x|x_{0},x_{1})\\ (x_{0},x_{1})\sim\pi^{*}\end{subarray}}\left[\left\lVert v_{\theta}(t,x)-u_{t}^{\circ}(x|x_{0},x_{1})\right\rVert^{2}+\lambda(t)^{2}\left\lVert s_{\theta}(t,x)-\nabla\log p_{t}(x|x_{0},x_{1})\right\rVert^{2}\right].$
A.1 Weighting Schedule $\lambda(t)$
Following Tong et al. [78], we apply a time-dependent weighting schedule $\lambda(t)$ to stabilize the score regression loss across time. This is essential because the conditional score $\nabla_{x}\log p_{t}(x\mid x_{0},x_{1})$ grows unbounded as $t\rightarrow 0$ or $t\rightarrow 1$, leading to imbalance in the loss. To address this, we adopt the formulation that standardizes the regression target to have unit variance. This is done by predicting the noise added in sampling $x_{t}$ from the linear interpolation $\mu_{t}=(1-t)x_{0}+tx_{1}$ under a Brownian bridge with variance $\sigma_{t}^{2}=\sigma^{2}t(1-t)$. This leads to:
@@ B.1 VAE Training
Hand Encoder. We use a PVCNN-based architecture [47] to process the input hand point cloud. The input to the encoder is a point cloud of size $N\times 3$ The PVCNN follows a hierarchical structure with point and voxel convolution blocks as described below:
@@ B.2 Score and Flow Model Architecture
Both the score network $s_{\theta}$ and the flow network $v_{\theta}$ in our pipeline share the same U-ViT-based backbone architecture [4]. The input to each model consists of a concatenation of latent hand representations, object shape encodings, and contact anchor tokens, following the conditioning design described in Section 4.3 of the main paper.
@@ Appendix C OT Cost Formulation and Computation
We form the convex hull of these wrenches to approximate the GWS, denoted as the grasp wrench hull (GWH). The distance between two grasps is then defined as the 1 minus the intersection-over-union (IoU) of their GWH volumes:
@@ D.2 Inference Time Comparison
We compare the inference time of our method against both a diffusion-based baseline and the optimization-based RobotFingerPrint [37] method. As shown in Table 8, our model supports fully amortized generation and remains efficient at inference time, taking approximately 0.8 seconds per grasp using 100 Euler–Maruyama steps. This is comparable to the diffusion baseline (0.5s) and significantly faster than RFP, which requires costly per-grasp optimization.
Table 8: Average inference time per grasp (amortized).
@@
Diffusion (100 steps) 0.5
Ours (100 steps) 0.8
RFP [37] 5.2

### v2

@@ 5.1 Experimental Setup
[…]
@@
Metrics. We assess performance using several quantitative metrics. First, we report the grasp success rate using IsaacGym [57], following the GenDexGrasp [46] evaluation protocol. A grasp is considered successful if it maintains a stable hold on the object under six perturbation trials, where external forces are applied along $\pm x$, $\pm y$, and $\pm z$ axes. We apply each force as a uniform acceleration of $0.5m/s^{2}$ for 60 simulation steps and measure whether the object translates more than 2 cm. A grasp passes if it withstands all six trials. Simulation parameters include a friction coefficient of 10 and an object density of 10,000. We use IsaacGym’s built-in positional controller to achieve the desired joint configurations. In addition to the success rate, we report the diversity of the generated grasps, computed as the average standard deviation across all revolute joints, as well as the functional similarity between source and translated grasps. The latter is measured using the full 6D IoU of GWH as defined in Eq. 8.
Baselines. We evaluate all four variants of our model, each trained using a distinct physics-based OT cost introduced in Sec. 4.2. As baselines, we compare against several methods. We include three recent grasp transfer approaches: RobotFingerPrint [41], which uses a unified gripper coordinate space; Dex-Retargeting [72], an optimization-based method that retargets key points on finger links; and the neural retargeting module from CrossDex [98]. Notably, both Dex-Retargeting and CrossDex require pre-defined, manually annotated link-to-link correspondences, which limit their generality for arbitrary hand morphologies. Finally, we compare against a generative baseline: a diffusion-based model trained with samples from Contact Map Similarity OT and conditioned on the source contact map.
@@ 5.2 Quality of Translated Grasps
Our method significantly outperforms all baselines in terms of grasp success rate. Our approach using the Jacobian-based OT cost and GWH-based variants achieves the highest mean success rates (67.45% and 66.34%, respectively), which suggests that both wrench- and control-aligned transport costs are particularly effective for grasp transfer. The IK-style optimization baselines, Dex-Retargeting [72] and CrossDex Retargeting Module [98], perform poorly. Their focus on select links often ignores collisions with other parts of the hand (e.g., the palm), leading to a high rate of physically invalid grasps. In contrast, our generative approach learns to produce configurations that respect the full-hand morphology. Qualitative results in Fig. 4 (left) further illustrate that our method generates stable grasps even from sub-optimal source poses where baselines fail.
Figure 4: (left) Qualitative results. The ‘Pose’, ‘Contact’, ‘GWH’, and ‘Jacobian’ columns show results from our method when trained using the respective OT cost functions from Sec. 4.2. Our method generates stable and consistent grasps across different OT variants, even when the source grasp is sub-optimal, where RFP [41] fails. (right) Failure Modes where our method struggles with thin-shell objects due to challenging geometry.
On the diversity metric, our contact-based variant achieves the highest mean diversity, producing a wide range of plausible grasps. This highlights that our methods can generate diverse grasp configurations while maintaining physical feasibility. Interestingly, other OT costs show lower diversity despite achieving high success rates, indicating that geometric alignment alone may lead to more conservative grasps. While the diffusion baseline is also highly diverse, our method achieves this while demonstrating superior performance on stability and functional alignment.
In addition to performance, our approach offers significant computational advantages over optimization-based baselines. Methods like RobotFingerPrint [41] require costly per-grasp iterative optimization, leading to an average inference time of over 5 seconds. In contrast, our generative method amortizes this cost through batch processing. With 100 integration steps, our model achieves an average inference time of approximately 0.8 seconds per grasp, comparable to the diffusion baseline and substantially faster than RobotFingerPrint. While lightweight methods like CrossDex [98] are faster due to their single-pass architecture, our approach provides a strong balance between computational efficiency and the generation of high-quality, physically robust grasps. A more detailed efficiency report is included in Appendix D.2.
@@ 5.3 Physical Alignment between Translated Grasps
To evaluate whether our translated grasps preserve the functional intent of the source grasp, we compute the 6D GWH IoU between the source and the translated target grasps. As shown in Table 1 (right block), our methods trained with the GWH-based and Contact-based OT costs achieve the highest mean IoU by a substantial margin (14.63% and 14.54%, respectively), significantly outperforming all baselines. This indicates that our approach more effectively preserves the object-level physical capabilities of the grasp, rather than simply mimicking coarse finger geometry. We provide a complementary analysis on the geometric alignment of contact points in Appendix D.3, which further supports these findings.
Taken together, our results reveal a key trade-off between the different physics-informed objectives. The Jacobian-based cost yields the highest success rate, suggesting it is most effective at capturing the kinematic relationships required for stability. However, the GWH and Contact costs achieve superior functional alignment, as measured by GWH IoU. This indicates that directly optimizing for force-exertion capabilities (GWH) or contact patterns (Contact) is more effective for transferring the grasp’s underlying mechanical intent, even if this sometimes results in a slightly lower stability score compared to a pure manipulability-based objective.
5.4 Ablation Study
@@ 6 Discussion
Limitations. Despite its strengths, our approach has some limitations. Most notably, as shown in Fig. 4 (right), the performance drops in scenarios involving thin-shell objects with ambiguous or minimal contact regions, where the valid grasp configuration space is largely confined. Additionally, we observe that the results for the Shadow hand are consistently lower than for the Allegro hand across multiple metrics, which aligns with prior findings in the MultiGripperGrasp dataset [12] due to its more complex kinematics and larger workspace. Furthermore, our current framework does not generalize to unseen hand morphologies at inference time. The VAE decoder implicitly learns the kinematic structure of the target hand, and would thus require retraining or fine-tuning on data from a new hand to generate valid grasps. Finally, these discrepancies suggest that improving dataset quality and coverage for specific hands may further enhance generalization and translation fidelity.
@@
Appendix A Background on Schrödinger Bridges and Simulation-Free Training
A.1 The Schrödinger Bridge Problem
The Schrödinger Bridge (SB) problem seeks a stochastic process $\mathbb{P}$ that evolves from a source distribution $q_{0}(x)$ to a target distribution $q_{1}(x)$ while being as "close" as possible to a reference process $\mathbb{Q}$ (typically Brownian motion). This is formally expressed as minimizing the Kullback-Leibler (KL) divergence between the path measures:
The optimal process $\mathbb{P}^{*}$ can be characterized as a mixture of Brownian bridges weighted by an entropically regularized optimal transport (OT) plan [21, 11]. As shown in Eq. 1 in the main text, this plan $\pi_{\varepsilon}^{*}$ finds a low-cost coupling between $q_{0}$ and $q_{1}$. The work of Föllmer [31] shows that if $\mathbb{Q}$ is Brownian motion, the SB is uniquely described by a mixture of Brownian bridges drawn according to $\pi_{\varepsilon}^{*}$. This insight is crucial, as it reduces a complex dynamic optimization problem to a more tractable computation over static distributions.
A.2 Learning Stochastic Dynamics with [SF]2M
The dynamics of the SB are modeled by the Itô SDE in Eq. 2. The marginal densities $p_{t}(x)$ induced by this SDE evolve according to the Fokker-Planck equation [73]:
where $\Delta p_{t}=\nabla\cdot(\nabla p_{t})$ is the Laplacian. The probability flow ODE in Eq. 3 shares the same marginals $p_{t}(x)$. To learn these dynamics from data without direct access to the marginals $p_{t}$, the [SF]2M framework [84] leverages the connection to entropic OT. By sampling a source-target pair $(x_{0},x_{1})$ from the entropic OT plan $\pi^{*}_{\varepsilon}$, one can construct a conditional Brownian bridge path between them. For a constant diffusion rate $g(t)=\sigma$, the conditional distribution $p_{t}(x\mid x_{0},x_{1})$ is a Gaussian $\mathcal{N}(x;(1-t)x_{0}+tx_{1},\sigma^{2}t(1-t)I)$. From this, we obtain closed-form expressions for the conditional score and probability flow drift, which serve as regression targets for the neural networks:
$\begin{split}u_{t}^{\circ}(x|x_{0},x_{1})&=\frac{1-2t}{t(1-t)}(x-((1-t)x_{0}+tx_{1}))+(x_{1}-x_{0}),\\ \nabla\log p_{t}(x|x_{0},x_{1})&=\frac{(1-t)x_{0}+tx_{1}-x}{\sigma^{2}t(1-t)}.\end{split}$ (12)
These targets are used in the loss function shown in Eq. 4.
A.3 Weighting Schedule $\lambda(t)$
Following Tong et al. [84], we apply a time-dependent weighting schedule $\lambda(t)$ to stabilize the score regression loss across time. This is essential because the conditional score $\nabla_{x}\log p_{t}(x\mid x_{0},x_{1})$ grows unbounded as $t\rightarrow 0$ or $t\rightarrow 1$, leading to an imbalance in the loss. To address this, we adopt the formulation that standardizes the regression target to have unit variance. This is done by predicting the noise added in sampling $x_{t}$ from the linear interpolation $\mu_{t}=(1-t)x_{0}+tx_{1}$ under a Brownian bridge with variance $\sigma_{t}^{2}=\sigma^{2}t(1-t)$. This leads to:
@@ B.1 VAE Training
Hand Encoder. We use a PVCNN-based architecture [51] to process the input hand point cloud. The input to the encoder is a point cloud of size $N\times 3$. The PVCNN follows a hierarchical structure with point and voxel convolution blocks as described below:
@@ B.2 Score and Flow Model Architecture
Both the score network $s_{\theta}$ and the flow network $v_{\theta}$ in our pipeline share the same U-ViT-based backbone architecture [5]. The input to each model consists of a concatenation of latent hand representations, object shape encodings, and contact anchor tokens, following the conditioning design described in Sec. 4.3 of the main paper.
@@ Appendix C OT Cost Formulation and Computation
We form the convex hull of these wrenches to approximate the GWS, denoted as the grasp wrench hull (GWH). The distance between two grasps is then defined as 1 minus the intersection-over-union (IoU) of their GWH volumes:
@@ D.2 Inference Time Comparison
We compare the amortized inference time of our method against all baselines in Table 8. While our framework is designed primarily as an offline planner, its inference speed is competitive.
As shown in the table, our method takes approximately 0.8 seconds to generate a grasp using 100 Euler-Maruyama steps. This is comparable to the diffusion baseline (0.5s). The iterative sampling process, where each of the 100 steps requires a full forward pass through our U-ViT model, is the primary factor for this timing. In contrast, lightweight methods like CrossDex are significantly faster, as they only require a single forward pass through a small MLP. Nevertheless, our approach is substantially faster than the optimization-based RobotFingerPrint, which requires over 5 seconds per grasp due to its costly per-instance optimization procedure.
Table 8: Average inference time per grasp (amortized, in seconds). Our method is comparable to the diffusion baseline and significantly faster than optimization-based approaches.
@@
[…]
CrossDex Retargeting [98] 1.1e-5 1.1e-5 1.1e-5 1.1e-5
Dex-Retargeting [72] 0.13 0.16 0.08 0.12
Diffusion (100 steps) 0.5 0.5 0.5 0.5
Ours (100 steps) 0.8 0.8 0.8 0.8
RobotFingerPrint [41] 5.2 5.8 5.1 5.37
D.3 Contact Alignment
In addition to the GWH IoU metric presented in the main text, we evaluate the geometric alignment of contact patterns using a Contact Alignment metric. This metric is defined as the squared Chamfer distance between the source and target contact maps on the object surface, where a lower value indicates better alignment. This serves as a useful proxy for how well the local interaction geometry of the grasp is preserved during translation.
[…]
RobotFingerPrint [41] 19.21 16.66 11.27 15.71
Dex-Retargeting [72] 9.53 12.92 6.68 9.71
CrossDex Retargeting [98] 9.85 14.52 7.06 10.48
Ours (Contact OT) 5.05 3.37 6.54 4.99
Ours (GWH OT) 10.71 7.26 10.19 9.39
D.4 VAE Ablation Study
To validate the design choices for our Variational Autoencoder (VAE), we conduct an ablation study comparing our full model against three alternatives:
  • PointNet VAE: A baseline inspired by [1] that reconstructs the hand point cloud via Chamfer distance and decodes hand parameters in a separate branch.
  • Ours w/o Mesh Loss: Our VAE trained only on the hand parameter reconstruction loss ($\mathcal{L}_{VAE}=\mathbb{E}[\left\lVert\hat{g}-g\right\rVert^{2}]$), without the differentiable FK layer and mesh vertex loss.
  • Ours w/o Param Loss: Our VAE trained only on the mesh vertex loss ($\mathcal{L}_{VAE}=\mathbb{E}[\left\lVert\hat{v}-\mathrm{FK}(g)\right\rVert^{2}]$), without the direct parameter reconstruction loss.
We evaluate reconstruction quality using the L2 error on the decoded hand parameters, tested on a held-out set of objects. As shown in Table 10, our proposed VAE architecture significantly outperforms all alternatives.
[…]
Method Allegro Shadow Human
PointNet VAE 0.1965 0.2001 0.2468
Ours w/o Mesh Loss 0.0967 0.0933 0.1270
Ours w/o Param Loss 0.0439 0.0441 0.0534
Ours (Full) 0.0335 0.0286 0.0523
Our analysis indicates that the PointNet VAE struggles because the Chamfer distance loss can be indiscriminative; it may achieve a low value even if the reconstructed pose is incorrect but has large overlapping regions with the ground truth. Ablating the mesh loss harms the learning of fine-grained geometric details, while ablating the parameter loss can lead to memorization issues, as the high-capacity latent space (dim=768) can overfit to the relatively low-dimensional hand configuration space (typically $\sim$20 DoFs).
D.5 Results on Additional Transfer Settings
[…]
RobotFingerPrint 30.27
Diffusion 33.96
Ours (Contact OT) 38.00
Ours (GWH OT) 41.49
[…]
RobotFingerPrint 0.196 0.180 0.182 0.186
Diffusion 0.283 0.298 0.207 0.263
Ours (Contact OT) 0.317 0.332 0.212 0.287
Ours (GWH OT) 0.310 0.327 0.178 0.272
[…]
RobotFingerPrint 3.46 4.11 3.51 3.69
Diffusion 10.41 11.82 8.78 10.34
Ours (Contact OT) 13.59 15.37 8.87 12.61
Ours (GWH OT) 11.76 13.30 9.16 11.41